# Estimating angular momenta of fission fragments from isomeric yield ratios using TALYS

### 22 Feb 2024 v1

@@ Abstract
To estimate the angular momenta of fission fragments based on the observed isomeric yield ratios in 25-MeV proton-induced fission of ${}^{238}$ and to compare these using Wilson’s model.
@@ Results
Data of 31 isomeric yield ratios in 25-MeV proton-induced fission of ${}^{238}$ were analysed. As a result, the average $J_{rms}$, equivalent to average angular momentum $J^{av}$, with uncertainties of 24 fission products, are presented. However, no conclusive result for the angular momentum could be obtained in seven cases. Considering the neutron emissions of the primary fission fragments, the average angular momentum as a function of the average mass number of the primary fission fragment is presented.
@@ I Introduction
Since the discovery of nuclear fission, considerable effort has been made to understand and describe this phenomenon. The scissioning nucleus, being a quantum many-body system, can not easily be described on the microscopic level, although considerable advancements have been made in the past decades [1, 2], and as a complement, macroscopic and phenomenological models are often used [3, 4, 5, 6, 7]. To optimize and validate these models, accurate experimental data from a broad range of fission reactions and observables are needed. One of the long-standing questions in fission dynamics is the large angular momenta of the fission fragments. Angular momentum of fission fragments have been derived from experimental observables via indirect methods [8, 9, 10, 11, 12, 13, 14, 15], and the results show that the root-mean-square angular momenta typically are around 6 - 7 $\hbar$. This is, in many cases, considerably higher than those of the fissioning nuclei, which raises the question of how these large angular momenta are generated in the fission process.
J. Wilson et al. [15] have proposed that the angular momentum is generated after scission from the collective motion of nucleons in the ruptured neck of the fissioning system. This conclusion is based on observed average angular momenta deduced from $\gamma$-ray spectroscopy, which seems to suggest that the angular momentum of a fission fragment is independent of that of the complementary fragment. Wilson et al. also propose a model for the angular momentum as a function of mass for fission products in the low and high mass peaks. However, there is no systematic study, and hence no model, for the angular momenta of fission fragments in the symmetric mass region.
@@
In early attempts, a statistical method based on the Rayleigh distribution [17, 18], used to describe the angular momentum distribution of the excited nucleus, was used to derive the angular momenta from measured isomeric yield ratios [8, 9, 12, 11]. The isomeric yield ratio (IYR), in general terms referring to the relative yield of different long-lived states of a fission product, is in this context more specifically defined as the relative yield of the high spin state to the total yield of the fission product.
@@
Other fission models, such as MCHF [4], CGMF [5], FIFRELIN [6] and FREYA [7], use the Hauser-Feshbach formalism [22] to describe the de-excitation of the fission fragments, linking the angular momenta to the IYRs. However, these models also use assumptions and model parameters to describe the path from compound nuclei to fission fragments, which are not validated with the nuclear data from proton-induced fission. In this study, GEF is only used to provide the initial states of the fission fragments, which will be tuned in the following calculations. Hence, GEF is good enough to be employed in this study.
In an alternative approach to link the angular momenta to the measured IYRs, the reaction code TALYS [23], which also uses the Hauser-Feshbach formalism for the deexcitation, has been used to calculate the yields of the isomers starting from different angular momentum distributions [24, 13, 14]. In these studies, a clear and strong correlation between the angular momentum and IYR has been observed. Based on that correlation, the average root-mean-square angular momentum $J_{rms}$, determining the angular momentum distribution, has been deduced. However, also in this approach, several assumptions and simplifications are made that might affect the results. For example, the contribution from each FF to the production of the isomer is not calculated. Without that information, it is too simple that three FFs or even fewer are considered in calculating the isomeric yield ratio. The calculated IYR from each FF is compared with the experimental value to deduce the $J_{rms}$, and the fission yields are not considered in calculating the average $J_{rms}$. Furthermore, the excitation energy of the excited FFs is either modelled using only the mean energy or as a Gaussian distribution. In both cases, the excitation energy is assumed to be independent of the angular momentum distribution. This means that when the value of $J_{rms}$ is increased to an extreme value, the population of the FFs might exceed the yrast line, which is nonphysical.
In order to have a more realistic description of the angular momentum and energy distributions of the FFs, TALYS has been coupled to GEF. In this approach, TALYS is used to calculate the yield of the isomers, starting with the properties of the fission fragments extracted from GEF simulations. The results were evaluated by comparing the derived IYR with measured values, and the conclusion was that the angular momenta of the FFs from GEF was underestimated [21].
The present work builds on the GEF + TALYS approach by developing a surrogate model for GEF, in which the $J_{rms}$ of the angular momentum distributions for the FFs can be varied without crossing the yrast line into an unphysical region. The goal is to estimate the most likely angular momenta of the fission fragments contributing to the yields of the isomers in the 31 measured IYRs from 25 MeV proton-induced fission of ${}^{238}$[21].
@@ II.1 Surrogate model
In the development of the model, GEF 2023/1.2 was used to provide the excitation energy (Ex) and angular momentum (J) distributions of the primary fission fragments in proton-induced fission of ${}^{238}$ at 25 MeV. Based on this GEF data, the energy vs angular momentum matrix (Ex-J) of the excited FFs was parameterised. These parameters could then be used to reconstruct the matrix and use it as input to TALYS, in order to derive the IYR. Thanks to the parameterisation, the parameter determining the angular momentum distribution of the input matrix could be tuned to reproduce the measured IYR.
@@ II.1.1 GEF: Ex-J matrix
Due to the high incident proton energy, up to three neutrons can be emitted before scission, resulting in four paths of fission in GEF. This pre-scission neutron emission changes the scissioning nucleus (SN) as well as its excitation energy. Hence, the properties of the primary fragments produced from different scissioning nuclei will be different.
As an example, Figure 1 shows the Ex-J matrix of the excited FF ${}^{133}$e* produced from the scissioning nucleus ${}^{237}$p* obtained from GEF, in color-filled contours. In the left and upper panels, the matrix is projected to the energy axis and angular momentum axis, respectively.
Figure 1: Ex-J matrix of the FF ${}^{133}$e* produced from the scissioning nucleus ${}^{237}$p* obtained from GEF, in color-filled contours. In the main panel, black points and error bars represent means and standard deviations of energies of the nuclei with a specific angular momentum. The red points show a fit to the average energy versus angular momentum. Red error bars are the standard deviation of energies, which are the same for all angular momentum. The left panel shows the excitation energy distribution from GEF in black points and a fitted normal distribution, N(19.6 MeV, 9.6 MeV), in red. The upper panel shows the angular momentum distribution from GEF in black points and a fit of the Rayleigh distribution, giving $J_{rms}=9.2~{}\hbar$.
In GEF, the angular momentum distribution of a particular FF follows the Rayleigh distribution [17, 18],
@@
The energy distribution, shown in black points in the left panel of Figure 1, is reasonably well described by a normal distribution N(19.6 MeV, 9.6 MeV), except for the energy above 30 MeV. Such a distribution is good enough to serve as a surrogate model.
If, instead, the data is analysed for one value of the angular momentum at a time, the resulting energy distributions are well described by normal distributions. The means and standard deviations of these distributions are presented as black points with error bars in the main panel of Figure 1. It can be noticed that the standard deviations of the individual distributions are approximately the same, and can hence be replaced by their average value $\bar{\sigma}_{E}$.
From Figure 1, it is observed that the mean energy for each angular momenta ($\bar{E}_{ex}$) correlates with the angular momenta. Following the parameterisation suggested by Al-Adili et al. [25] the relations ship can be described by
@@
in which $C$ and $\bar{E}_{offset}$ are free parameters. In the case of ${}^{133}$e, the correlation is written as $\bar{E}_{ex}=0.037\cdot J(J+1)+16.0$ MeV.
To summarise, a set of four parameters: $J_{rms}$, $C$, $\bar{E}_{offset}$, and $\bar{\sigma}_{E}$, are needed to describe the Ex-J matrix of a single FF from one particular scissioning nucleus. Considering the four possible scissioning nucleus in the fission of ${}^{238}$ at 25 MeV, a total of 20 parameters, also including the four probabilities for each scissioning nuclei, are needed to describe the Ex-J matrix. As an example, the values of 20 parameters describing the four Ex-J matrices for ${}^{133}$e* are listed in Table 1.
@@
($\hbar$) MeV MeV
@@ II.1.2 Reconstruction of the Ex-J matrix
Figure 2 shows the reconstructed Ex-J matrix of the FF ${}^{133}$e*, which is a weighted sum of four matrices. Each matrix is reconstructed using a set of three parameters: $C$, $\bar{E}_{offset}$, and $\bar{\sigma}_{E}$ from Table 1, and $J_{rms}$ = 9.0 $\hbar$. Projections of the matrix to the energy and angular momentum axes are presented in red curves in the left and upper panels, respectively. The corresponding projections of the Ex-J matrix from GEF are presented as black lines. The projected angular momentum of the matrix is a good approximation of the GEF data. This means that the single values of $J_{rms}$ could be used to replace the four values in Table 1, without serious loss of information.
Figure 2: Reconstructed Ex-J matrix of the primary FF ${}^{133}$e, for all four scissioning nuclei. The upper and left panels show the angular momentum distribution and the excitation energy distribution from GEF in black dots and corresponding projections of the reconstructed matrix as red curves.
To study how the angular momentum affects the de-excitation of the FF $J_{rms}$ is varied, and the resulting matrices are fed to TALYS. As an example, Figure 3 shows a reconstructed matrix with a $J_{rms}$ of $15.0~{}\hbar$. In the upper panel of Figure 3 it can be seen that the reconstructed angular momentum distribution changes as expected. At the same time, the energy distribution is also significantly changed (left panel of Figure 3) due to the dependency on the angular momentum described by Eq 3.
Figure 3: Reconstructed Ex-J matrix with $J_{rms}=15.0~{}\hbar$ for the FF ${}^{133}$e* based on the output from GEF. The upper and left panels show the angular momentum distribution and the excitation energy distribution from GEF in black dots and corresponding projections of the reconstructed matrix by red curves.
@@ II.2 De-excitation with TALYS
TALYS [23] is a general nuclear reaction code commonly used in data evaluation. In this study, the latest version, TALYS 1.96 [26], is used to calculate the de-excitation of excited FFs generated by the surrogate model.
@@
To predict the level densities in the continuum above the highest discrete energy level, and to supplement the discrete levels whenever the level scheme is incomplete, the Back-shifted Fermi gas Model (BFM) [27] was used as the default option, In earlier studies, impacts of the choice of level density model on the calculated IYRs have been studied [24, 13]. Different level density models were therefore tested, and the BFM model showed better performance than the other models in most cases. However, for cadmium, indium and antimony, the microscopic level densities calculated by S. Goriely showed a better performance and were therefore used instead.
As pointed out by V. Rakopoulos [14], and also observed in the current work, the knowledge of the discrete level scheme can have a significant impact on the calculation of the IYRs in TALYS. Hence, the level schemes used in the calculations were updated with the latest experimental data retrieved from the RIPL-3 database [28]. In all calculations, TALYS was set up to use all experimentally known levels up to the fortieth level, which is the upper limit in TALYS.
Provided with an input file, TALYS calculates the de-excitation of the excited nucleus and prints the results to an output file. From this output, the population of the long-lived high-spin ($P_{hs}$) and low-spin ($P_{ls}$) states are extracted. For the FF ${}^{133}$e*, three of the products after de-excitation, for which experimental data of the IYRs are available for comparison, are ${}^{129,131,133}$e. Figure 4 shows the population of these three products as a function of $J_{rms}$ in colored solid curves, while the black curve shows the total population of the three products. Considering that the population is normalized to the population of the initial fragment, about half of the FFs ${}^{133}$$*{}^{*}$ de-excites to either of these three products. Hence, observations from these three FPs provide an important part of the information necessary to estimate the $J_{rms}$ of the FF ${}^{133}$e*.
Figure 4: Populations of the fission products from the de-excitation of ${}^{133}$e* as a function of $J_{rms}$.
@@
Furthermore, it can be observed in Figure 4 how the summed population of two states of ${}^{129}$e increases with $J_{rms}$, while the population of ${}^{133}$e decreases. The population of ${}^{131}$e increases up to about $10.6~{}\hbar$, and then starts to decrease. Such dependencies are not surprising, considering that the excitation energy of the fragment increases with angular momentum (see Eq. 3), leading to larger neutron emission.

### 18 Mar 2024 v3

@@ Abstract
To estimate the angular momenta of fission fragments based on the observed isomeric yield ratios in 25-MeV proton-induced fission of ${}^{238}$.
@@ Results
Data of 31 isomeric yield ratios in 25-MeV proton-induced fission of ${}^{238}$ were analysed. From the analysis, the average $J_{rms}$, equivalent to average angular momentum $J^{av}$, with uncertainties are obtained for 24 fission products, while in seven cases no conclusive result for the angular momentum could be obtained. Furthermore, considering the neutron emissions of the primary fission fragments, the average angular momentum as a function of the average mass number of the primary fission fragment was estimated.
@@ I Introduction
Since the discovery of nuclear fission, considerable effort has been made to understand and describe this phenomenon. The scissioning nucleus, being a quantum many-body system, can not easily be described on the microscopic level, although considerable advancements have been made in the past decades [1, 2]. As a complement, macroscopic and phenomenological models are often used [3, 4, 5, 6, 7]. To optimize and validate these models, accurate experimental data from a broad range of fission reactions and observables are needed. One of the long-standing questions in fission dynamics is the observation of large angular momenta of the fission fragments. Angular momenta of fission fragments are derived from experimental observables via indirect methods [8, 9, 10, 11, 12, 13, 14, 15], and the results show that the root-mean-square angular momenta typically are around 6 - 7 $\hbar$. This is, in many cases, considerably higher than those of the fissioning nuclei, which raises the question of how these large angular momenta are generated in the fission process.
In a recent article, J. Wilson et al. [15] have proposed that the angular momenta are generated after scission, from the collective motion of nucleons in the ruptured neck of the fissioning system. This conclusion is based on fission fragments’ angular momenta estimated from $\gamma$-ray spectroscopy, which seems to suggest that the angular momentum of a fission fragment is independent of that of the complementary fragment. Wilson et al. also propose a parameterization of the angular momentum as a function of mass,d for fission products in the low and high mass peaks.
However, competing explanations for the angular momenta observed by Wilson et al. exist. For example, Randrup et al. [16, 17] claim that the angular momentum is generated before scission. This would mean that the angular momenta of a pair of complementary fission fragments are correlated, but that the correlation may vanish during scission. Another explanation for the lack of correlation is put forward by Stetcu et al.. They estimate the angular momentum removal by prompt neutrons and $\gamma$-rays to be in the order of 3.5 - 5 $\hbar$ [18]. In their opinion, such a wide distribution of angular momentum removal can hide any underlying correlations of fission fragments’ angular momenta. Bulgac et al. uses a microscopic model to calculate the angular momenta of fission fragments, and suggest that it is the collective spin modes of the fission fragments that dominate the angular momentum generation [2].
So far, there is no systematic study, and hence no model, for the angular momenta of fission fragments in the symmetric mass region. Hence, more data on the angular momenta of fission fragments, in particular in the symmetric region, are needed.
@@
In early attempts, a statistical method based on the Rayleigh distribution [20, 21], used to describe the angular momentum distribution of the excited nucleus, was used to derive the angular momenta from measured isomeric yield ratios [8, 9, 12, 11]. The isomeric yield ratio (IYR), in general terms referring to the relative yield of different long-lived states of a FP, is in this context more specifically defined as the relative yield of the high spin state to the total yield of the FP.
@@
Other fission models, such as MCHF [4], CGMF [5], and FIFRELIN [6], use the Hauser-Feshbach formalism [25] to describe the de-excitation of the FFs, linking the angular momenta to the IYRs. However, these models also use assumptions and model parameters to describe the path from compound nuclei to FFs, which are generally not validated against nuclear data from proton-induced fission.
In an alternative approach to link the angular momenta to the measured IYRs, the reaction code TALYS [26], which also uses the Hauser-Feshbach formalism to model the deexcitation, has been used to calculate the yields of the isomers starting from different angular momentum distributions [27, 13, 14]. In these studies, a clear and strong correlation between the angular momentum and IYR has been observed. Based on that correlation, the average root-mean-square angular momentum $J_{rms}$, determining the angular momentum distribution, has been deduced. However, also in this approach, several assumptions and simplifications are made that might affect the results. For example, the contribution from each FF to the production of the isomer is not calculated, and only three FFs, or even fewer, are considered to contribute to the measured isomeric yield ratio. Furthermore, the calculated IYR from each FF is compared with the experimental value to deduce the $J_{rms}$, and the fission yields are not considered in calculating the average $J_{rms}$. Also, the excitation energy of the excited FFs is either modelled using only the mean energy or as a Gaussian distribution. In both cases, the excitation energy is assumed to be independent of the angular momentum distribution. This means that when the value of $J_{rms}$ is increased to an extreme value, the population of the FFs might exceed the yrast line, which is unphysical.
In order to have a more realistic description of the angular momentum and energy distributions of the FFs, TALYS has been coupled to GEF. In this approach, TALYS is used to calculate the yield of the isomers, starting with the properties of the FFs extracted from GEF simulations. The results were evaluated by comparing the derived IYR with measured values, and the conclusion was that the angular momenta of the FFs from GEF was underestimated [24].
The present work builds on the GEF + TALYS approach by developing a surrogate model for GEF, in which the $J_{rms}$ of the angular momentum distributions for the FFs can be varied without crossing the yrast line into an unphysical region. The goal is to estimate the most likely angular momenta of the FFs contributing to the yields of the isomers in the 31 measured IYRs from 25 MeV proton-induced fission of ${}^{238}$[24].
@@ II.1 Surrogate model
In this study, GEF is used to provide an initial guess of the states of the FFs. The properties of these states are then tuned using a surrogate model, and used as starting conditions of TALYS calculations.
In the development of the model, GEF 2023/1.2 was used to provide the excitation energy (Ex) and angular momentum (J) distributions of the primary FFs in proton-induced fission of ${}^{238}$ at 25 MeV. Based on this GEF data, the energy vs angular momentum matrix (Ex-J) of the excited FFs was parameterised. These parameters could then be used to reconstruct the matrix and use it as input to TALYS, in order to derive the IYR. Thanks to the parameterisation, the parameter determining the angular momentum distribution of the input matrix could be tuned to reproduce the measured IYR.
@@ II.1.1 GEF: Ex-J matrix
Due to the high incident proton energy up to three neutrons can be emitted before scission, resulting in four paths to fission according to GEF. This pre-scission neutron emission changes the scissioning nucleus (SN) as well as its excitation energy. Hence, the properties of the primary fragments produced from different scissioning nuclei will be different.
As an example, Figure 1 shows the Ex-J matrix of the excited FF ${}^{133}$e* produced from the scissioning nucleus ${}^{237}$p*, as obtained from GEF, in color-filled contours. In the left and upper panels, the matrix is projected to the energy axis and angular momentum axis, respectively.
Figure 1: Ex-J matrix of the FF ${}^{133}$e* produced from the scissioning nucleus ${}^{237}$p*, as obtained from GEF, in color-filled contours. In the main panel, black points and bars represent the means and standard deviations of energy distributions for nuclei with a specific angular momentum. The red points show a fit to the average energy versus angular momentum. Red bars are the standard deviation of the fitted energies, assumed to be the same for all angular momenta. The left panel shows the excitation energy distribution from GEF in black points and a fitted normal distribution, N(19.6 MeV, 9.6 MeV), in red. The upper panel shows the angular momentum distribution from GEF in black points and a fit of the Rayleigh distribution, giving $J_{rms}=9.2~{}\hbar$.
In GEF, the angular momentum distribution of a particular FF is described by the Rayleigh distribution [20, 21],
@@
The energy distribution, shown in black points in the left panel of Figure 1, is reasonably well described by a normal distribution N(19.6 MeV, 9.6 MeV), except for energies above 30 MeV. Such a distribution is good enough to serve as a surrogate model.
If, instead, the data is analysed for one value of the angular momentum at a time, the resulting energy distributions are well described by normal distributions. The means and standard deviations of these distributions are presented as black points with error bars in the main panel of Figure 1. It can be noticed that the standard deviations of the individual distributions are approximately the same, and are hence replaced by their average value $\bar{\sigma}_{E}$.
From Figure 1, it is observed that the mean energy for each angular momenta ($\bar{E}_{ex}$) correlates with the angular momentum. Following the parameterisation suggested by Al-Adili et al. [28] the relations ship can be described by
@@
in which $C$ and $\bar{E}_{offset}$ are free parameters. In the case of ${}^{133}$e, the best fit to the data is $C~{}=~{}0.037$ MeV/2superscriptPlanck-constant-over-$\hbar^{2}$ and $\bar{E}_{offset}~{}=~{}16.0$ MeV.
To summarise, a set of four parameters: $J_{rms}$, $C$, $\bar{E}_{offset}$, and $\bar{\sigma}_{E}$, are needed to describe the Ex-J matrix of a single FF from one particular scissioning nucleus. Considering the four possible scissioning nuclei in the proton induced fission of ${}^{238}$ at 25 MeV, a total of 20 parameters, adding the four probabilities for each scissioning nuclei, are needed to describe the Ex-J matrix. As an example, the values of the 20 parameters describing the four Ex-J matrices for ${}^{133}$e* are listed in Table 1.
@@
$\hbar$ MeV/2superscriptPlanck-constant-over-$\hbar^{2}$ MeV MeV
@@ II.1.2 Reconstruction of the Ex-J matrix
Figure 2 shows the reconstructed Ex-J matrix of the FF ${}^{133}$e*, which is a weighted sum of four matrices. Each matrix is reconstructed using a set of three parameters: $C$, $\bar{E}_{offset}$, and $\bar{\sigma}_{E}$ from Table 1, and $J_{rms}$ = 9.0 $\hbar$. Projections of the matrix to the energy and angular momentum axes are presented in red curves in the left and upper panels, respectively. The corresponding projections of the Ex-J matrix from GEF are presented as black lines. The projected angular momentum of the matrix is a good approximation of the GEF data. This means that the single values of $J_{rms}$ could be used to replace the four values in Table 1, without serious loss of accuracy.
Figure 2: Reconstructed Ex-J matrix of the primary FF ${}^{133}$e, for all four scissioning nuclei. The upper and left panels show the angular momentum distribution and the excitation energy distribution from GEF in black dots and corresponding projections of the reconstructed matrix as red curves. The black dashed line, and the red dotted line represent the average excitation energy from GEF and the reconstructed matrix, respectivly.
To study how the angular momentum affects the de-excitation of the FF, $J_{rms}$ is varied and the resulting matrices are fed to TALYS. As an example, Figure 3 shows a reconstructed matrix for the FF ${}^{133}$e* with a $J_{rms}$ of $15.0~{}\hbar$. In the upper panel of Figure 3 it can be seen that the reconstructed angular momentum distribution changes as expected. At the same time, the energy distribution is also significantly changed (left panel of Figure 3) due to the dependency on the angular momentum described by Eq 3.
Figure 3: Reconstructed Ex-J matrix with $J_{rms}=15.0~{}\hbar$ for the FF ${}^{133}$e* based on the output from GEF. The upper and left panels show the angular momentum distribution and the excitation energy distribution from GEF in black dots and corresponding projections of the reconstructed matrix by red curves. The black dashed line, and the red dotted line represent the average excitation energy from GEF and the reconstructed matrix.
@@ II.2 De-excitation with TALYS
TALYS [26] is a general nuclear reaction code commonly used in data evaluation. In this study, TALYS 1.96 [29], is used to calculate the de-excitation of excited FFs generated by the surrogate model.
@@
To predict the level densities in the continuum above the largest known discrete energy level, and to supplement the discrete levels whenever the level scheme is incomplete, the Back-shifted Fermi gas Model (BFM) [30] was used as the default option. In earlier studies, impacts of the choice of level density model on the calculated IYRs have been observed [27, 13]. Different level density models were therefore tested, and it was concluded that the BFM model led to better agreement with the experimental observations than the other models in most cases. However, for cadmium, indium and antimony, the microscopic level densities calculated by S. Goriely led to better agreement with the experimental ratios and were therefore used instead.
As pointed out by V. Rakopoulos et al. [14], and also observed in the current work, the knowledge of the discrete level scheme can have a significant impact on the calculation of the IYRs in TALYS. Hence, the level schemes used in the calculations were updated with the latest experimental data retrieved from the RIPL-3 database [31]. In all calculations, TALYS was set up to use all experimentally known levels up to the fortieth level, which is the upper limit in TALYS.
Provided with an input file, TALYS calculates the de-excitation of the excited nucleus and prints the results to an output file. From this output, the population of the long-lived high-spin ($P_{hs}$) and low-spin ($P_{ls}$) states were extracted. For the FF ${}^{133}$e*, three of the products after de-excitation, for which experimental data of the IYRs are available for comparison, are ${}^{129,131,133}$e. Figure 4 shows the population of these three products as a function of $J_{rms}$ in colored solid curves, while the black curve shows the total population of the three products. Considering that the population is normalized to the population of the initial fragment, about half of the FFs ${}^{133}$$*{}^{*}$ de-excites to either of these three products. Hence, observations from these three FPs provide an important part of the information necessary to estimate the $J_{rms}$ of the FF ${}^{133}$e*.
Figure 4: Populations of the FPs from the de-excitation of FF ${}^{133}$e* as a function of $J_{rms}$.
@@
Furthermore, it can be observed in Figure 4 how the summed population of two states of ${}^{129}$e increases with $J_{rms}$, while the population of ${}^{133}$e decreases. The population of ${}^{131}$e increases up to about $10.6~{}\hbar$, and then starts to decrease. Such dependencies are not surprising, considering that the excitation energy of the fragment increases with angular momentum (see Eq. (3)), leading to larger neutron emission.